OpenStudy (anonymous):

What is the wavelength, in meters, of an alpha particle with a kinetic energy of 8.0x10^-13 J. The mass of an alpha particle= 4.00150 amu. Given 1 amu= 1.67x10^-27 kg. (K.E.= 1/2mv^2)

OpenStudy (masumanwar):

wave particle duality principles can be applied as hnu=mc^2=E

OpenStudy (anonymous):

I don't know how to use that formula

OpenStudy (michele_laino):

for example the momentum \(p\) of the alpha particle, is: \[\Large p = \sqrt {2 \cdot m \cdot KE} \] where \(KE\) is its kinetic energy, and \(m\) is its mass, namely: \[\Large \begin{gathered} KE = 8 \cdot {10^{ - 13}}joules \hfill \\ \\ m = 4.0015 \times {1.67\cdot 10^{ - 2t}} = ...?Kgr \hfill \\ \end{gathered} \] therefore, the requested wavelength \(\lambda\), is given by means of the formula of De Broglie: \[\Large \lambda = \frac{h}{p}\] where, \(h\) is the Planck's constant: \[\Large h = 6.62 \cdot {10^{ - 34}}joules \times \sec \]

OpenStudy (michele_laino):

oops.. it is: \[\Large m = 4.0015 \times 1.67 \cdot {10^{ - 27}} = ...?Kg\]